OpenStudy (anonymous):

sec theta = 8/3. theta is an acute angle, use a right trianlge to find the value of tan theta

OpenStudy (anonymous):

\[\sqrt{55}/3 ^{?}\]

OpenStudy (zzr0ck3r):

sec = 8/3 then cos = 3/8 so cah says 3=adjacent and 8 = hypotenuse so a^2+3^2=64 a=sqrt(55) so sin=sqrt(55)/8 so tan = (sqrt(55)/8)/(3/8)=sqrt(55)/3

OpenStudy (anonymous):

yeah i realized. thanks though